User Tools

Site Tools


handover:r1766

R1766

Notes

eremote still doesn't appear to work on ops8 OR ops4, the backup control doesn't appear to connect either (I could be wrong, I'm not very familiar with it). So the current way to view/monitor the experiment is to have the VNC to the pcfs open and have the logs going. Then you ssh into the pcfs from ops8 and find the 'stream_log' script.

It is currently located in /home/observer/deh/opsutils/ so cd to that dir and type (e.g for hb):

./stream_log hb

This will begin streaming the log from pcfs to /tmp/hb.log. Then from the log monitor, start monitoring that file (/tmp/hb.log). Seems to be best to do a clkoff first.

Because e-remote control is not working, checklists are unavailable. Will be regularly running through the checklist points and noting anything out of the ordinary below.

Hobart 12m

Disk vsn: BKG+0142/4000 Start position: 3.269 GBs

  • Autocorrelation spectrum at Hobart is a mess. Reconfigured dbbc and still no change. I'll see what happens after the experiment begins and if it fixes itself (Lucas). Update: after first scan it was fine. Must have been the bug in the plotting program.
  • Experiment started ok (Lucas).
  • 0042 - Ran fmset to correct jumps in reported pps_delay through the DBBC. (JMc)

Katherine

Disk vsn: HOB+0094/160000 Start position: 9051.643 GBs

  • Getting error 'computer to acu time difference exceeded 0.25 seconds, see value above' every second. All the delays are fine and the beep is extremely annoying. Wiki says error is meaningless (probably network issues?) but I cannot turn the beep off (Lucas).
  • Experiment started ok (Lucas).
  • 2200 - Rebooted drives in HMI to clear the ACU time difference problem. (JMc)

Yarragadee

Disk vsn: HOB0104/16000 Start position: 7877.132 GBs

  • Experiment started ok (Lucas).
  • 05:34 Wind stow, missed scan 320-0535b (Bryn)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/r1766.txt · Last modified: 2016/11/15 07:39 by Ross Turner